At 273.15 K and under a pressure of 10.1325 Mpa, the compressibility factor of Ois 0.927. Calculate the mass of Onecessary to fill a gas cylinder of 100 dmcapacity under the given conditions.

 

Grade:11

1 Answers

Akash
5 years ago
Z=PV/nRT
n=PV/ZRT
n=10.1325×10'(6)×10'(-6)×9.8×100/0.927/8.314/273.15
n=4.716
Mass= 4.716×32
Mass=151.9kg